A quality control analyst mixes 30.0 ml of 0.255M BaCl2 with excess AgNO3 causing Angel to precipitate and subsequently isolated 2.01 g of AgCl. Show a balanced equation for the reaction and calculate the percent yield of the reaction.
1
Expert's answer
2021-08-25T06:39:03-0400

BaCl2 + 2AgNO3 =2AgCl + Ba(NO3)2

According to the equation, n (AgNO3) = n (AgCl) = 2 x n (BaCl2)

CM = n / V

n = CM x V

n (BaCl2) = 0.255 x 0.03 = 0.008 mol

n (AgCl) = 0.008 x 2 = 0.016 mol

n = m / M

M (AgCl) = 143.32 g/mol

m (AgCl) = n x M = 0.016 x 143.32 = 2.29 g

% = 2.01/2.29 x 100 = 87.8%